Two rugby players run towards each other and have an inelastic collision. Given the mass 70 kg and 60 kg respectively, and velocity 3 ms-1 and 4 ms-1. Calculate the final velocity and its direction of the respective players. Discuss how a rugby player’s bone develops if continuously performing the weight bearing exercise using Wolff’s Law and bone remodeling theory.  


1
Expert's answer
2022-02-08T10:14:58-0500

Answer

inelastic collision so coefficient of restitution e=0

Mass of players

"m_1=70kg\\m_2=60kg"

Initial velocity

"u_1=3m/s\\u_2=4m/s"

Initial momentum

"P_i=m_1u_1+m_2u_2\\=-210+240\\=30kg-m/s"

So final velocities of players as below

"v_1=\frac{P_i}{(m_1+m_2) }\\=\frac{30}{(70+60) }\\=0.23m/s"

And also

"v_2=\frac{P_i}{(m_1+m_2) }\\=\frac{30}{(70+60) }\\=0.23m/s"

In this process momentum remains constant.